OpenStudy (anonymous):

I'm needing help trying to answer 3 problems, can anyone do a videochat with me on this stupid problems

OpenStudy (anonymous):

I can maybe help you through twiddla by using graphs and stuff. But first, let me see the question(s) so I can tell you if I'm qualified to help.

OpenStudy (saifoo.khan):

this site is full of Helpers! :)

OpenStudy (anonymous):

\[\sqrt{2x+4}=\sqrt{x+3}+1\]

OpenStudy (anonymous):

Hmm... I do think I can do this. Alright.

OpenStudy (saifoo.khan):

I can do this too!

OpenStudy (anonymous):

i gott stuck when you square the other side......

OpenStudy (anonymous):

Let's do it, then. :P http://www.twiddla.com/584304

OpenStudy (saifoo.khan):

the square of 1 is 1.

OpenStudy (vijay):

x = 6

myininaya (myininaya):

\[\sqrt{2x+4}=\sqrt{x+3}+1\] \[(\sqrt{2x+4})^2=(\sqrt{x+3}+1)^2\] \[2x+4=(\sqrt{x+3}+1)(\sqrt{x+3}+1)\] \[2x+4=(x+3)+2\sqrt{x+3}+1\] \[2x-x+4-3-1=2\sqrt{x+3}\] \[3x=2\sqrt{x+3}\] \[(3x)^2=(\sqrt{x+3})^2\] 9x^2=x+3

myininaya (myininaya):

oops my 2 disappeared!

OpenStudy (vijay):

nope, itz wrong myininaya

myininaya (myininaya):

\[9x^2=4(x+3)\]

myininaya (myininaya):

yes i know my i said my 2 disappeared lol

OpenStudy (vijay):

\[x ^{2} - 4x - 12 = 0\]

myininaya (myininaya):

ok you are right and i also did 2x+x instead of 2x-x you are right

OpenStudy (anonymous):

Which factors to (x-6)(x+2)

OpenStudy (anonymous):

We got it all solved in Twiddla. Lol. Great work, dude.

OpenStudy (phi):

Is there a convention about only taking the positive root? I.e. do we include the x= -2 as a solution?

OpenStudy (anonymous):

@Phi: You really go to MIT?

myininaya (myininaya):

-2 is not a solution since left hand side would not = right hand side sqrt{2(-2)+4}=0 sqrt{-2+3}+1=1+1=2

OpenStudy (anonymous):

:O Weird! How did I not catch that when I checked? Thanks, myininaya.

OpenStudy (phi):

OK, but in the second equation \[\sqrt{-2+3}= \sqrt{1}= \pm1\] So, if we choose -1 we would get -1+1= 0 But I think people only use the positive root??

myininaya (myininaya):

use only positive the symbol means principal square root (only positive outcome)